# Generating Verifiable Chain of Thoughts from Exection-Traces

### v2

@@
Abstract.
Teaching language models to reason about code execution remains a fundamental challenge. While Chain-of-Thought (CoT) prompting has shown promise, current synthetic training data suffers from a critical weakness: the reasoning steps are often plausible-sounding explanations generated by teacher models, not verifiable accounts of what the code actually does. This creates a troubling failure mode where models learn to mimic superficially convincing but logically flawed reasoning patterns.
We address this by grounding CoT generation directly in program execution traces. Our pipeline instruments code to capture its dynamic behavior, then narrates these execution traces into natural language and factually-grounded rationales that are verifiable by design. This execution-grounded approach ensures every reasoning step reflects what the program computes, eliminating logical hallucinations at the source. We evaluate our method on code reasoning tasks, code generation and explanation tasks from HumanEval. Models trained on our bi-directional trace-grounded data achieve substantial improvements on reasoning tasks, with gains of up to 30 points on output prediction and 28 points on input prediction over base models, alongside competitive explanation and code generation performance. https://github.com/shailja-thakur/Verified-Code-CoT
Chain-of-Thought, Code Reasoning, Execution-Traces, Synthetic Data Generation, Program Semantics, Large Language Models
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Software and its engineering General programming languagesccs: Computing methodologies Natural language processing
1. Introduction
Recent advances in large language models have enabled code assistants that can generate, explain, and debug programs through natural language interaction. These systems—exemplified by GitHub Copilot, Claude Code, and similar tools—have demonstrated remarkable syntactic fluency. However, they struggle with tasks requiring deep reasoning about program behavior: debugging requires understanding how state evolves, program comprehension demands tracing execution paths, and code refactoring needs verifying semantic preservation. These reasoning tasks are fundamental to software engineering practice, yet current models often produce plausible but incorrect explanations of program behavior.
From a programming languages perspective, this reveals a deeper challenge: while LLMs are powerful static models of source code, they fundamentally lack a connection to the dynamic behavior of programs—how state evolves during execution. This gap limits their utility to that of sophisticated auto-completes rather than true reasoning partners for tasks like debugging and program analysis, which require a faithful understanding of a program’s operational semantics.
To bridge this gap, the research community has embraced fine-tuning models on Chain-of-Thought (CoT) data, which provides explicit, step-by-step rationales to guide the model’s reasoning process. However, the methods used to generate this data for code are often unsound. A powerful teacher model generates the rationale, but these reasoning steps are not tethered to a program’s actual execution trace. The result is a dataset of logical hallucinations—plausible but factually incorrect accounts of the program’s logic.
This challenge is especially acute in code reasoning, where small logical errors cascade into flawed understanding. Consider a model trained on a rationale claiming ”variable count is incremented in the loop”—when in fact, count is decremented. Such errors propagate during training, teaching the model to follow plausible but incorrect logical patterns. This lack of guaranteed soundness is a critical barrier for high-stakes applications. For debugging, a rationale that misrepresents variable state is misleading. For iterative self-refinement, feeding a model its own flawed reasoning causes compounding errors. For agentic workflows, a plan based on unsound understanding of a previous step can derail an entire task. Figure 1 illustrates this critical problem concretely.
Figure 1. Comparison of hallucinated vs trace-grounded CoT. Our approach translates pysnooper (pysnooper) execution trace (showing line numbers, variable states, state transitions) into natural language reasoning. Color mapping shows trace sections: yellow=init, blue=iterations, green=return. LLM hallucinates wrong comparison (arr[2] vs arr[1] instead of arr[3]) and wrong search direction.
Existing work uses execution in two ways. Benchmarks like HumanEval (chen2021evaluating) and methods like SemCoder (ding2024semcodertrainingcodelanguage) validate that the final program output is correct, but do not verify the logical integrity of intermediate reasoning steps. Formal verification methods can provide mathematical guarantees but are computationally prohibitive at scale. A critical gap exists: there is no practical method for validating each step in the reasoning chain—what we term rationale-step verification—that scales to large-scale training data generation. Additionally, existing pipelines lack fine-grained control over problem complexity and diversity.
Furthermore, effective code reasoning requires both forward inference (predicting outputs from inputs) and backward reasoning (understanding how outputs arose from inputs) (li2025codeio). This bi-directional capability is essential for debugging, code comprehension, and program analysis. However, current synthetic data generation methods do not systematically address this requirement, leading to models with asymmetric reasoning capabilities.
In this work, we bridge these gaps by introducing a new paradigm for data synthesis that achieves rationale-step verification at scale. We have developed a complete synthesis pipeline that generates CoT rationales by directly translating program execution traces into natural language. Specifically, we instrument a program to capture its dynamic execution state—recording variable values after each statement, control flow decisions, and state transitions derived from program semantic operations. This trace serves as an ground truth from which rationales are derived. Each step in the resulting reasoning chain can be verified against the trace to confirm: (1) variable values match recorded runtime values, (2) state transitions correctly reflect how variables changed between statements, and (3) control flow accurately describes which branches executed. This ensures rationales are verifiable by design, eliminating the logical hallucinations that plague other methods.
@@ 1. Introduction
Figure 2. An overview of our three-stage data synthesis pipeline. Stage A (Concept Sourcing & Synthesis): Generates candidate concepts, code, and tests from raw documents. Stage B (Verification & Clustering): Uses our execution-based Dual Agreement algorithm to identify and rank the highest-quality solution-test pairs. Stage C (CoT Generation): Uses the verified artifacts and their execution traces to produce the final, bi-directional conversational data, complete with trace-grounded rationales and feedback.
@@
Table 1. Comparison of features in related works on code reasoning data and methodologies. Our work is the first to systematically combine execution-grounded verification with bi-directional CoT generation at scale.
Work / Method Execution-Trace Natural Language Bi-Directional Verifies Scalable Data
Grounded CoT CoT CoT Output Synthesis
REVTHINK (chen2024reverse)
TRACED (ding2024traced)
CodeI/O (li2025codeio)
rStar-Coder (liu2025rstarcoderscalingcompetitivecode)
Jung et al. (jung2025codeexecutiongroundedsupervision)
SemCoder (ding2024semcodertrainingcodelanguage)
@@
Furthermore, our concept-first pipeline provides fine-grained control over data quality. Rather than relying on existing code corpora, it synthesizes problems from abstract programming concepts (e.g., ”pass-by-reference semantics,” ”dynamic programming”) extracted from technical literature. This allows us to systematically control problem difficulty and ensure diverse coverage of algorithmic and mathematical domains. Combined with bi-directional trace generation, this provides a complete foundation for training models on both forward reasoning (predicting outputs) and backward reasoning (diagnosing how outputs arose), with all rationales grounded in verifiable execution.
Our contributions are fivefold:
  1. (1)
    Chain of Thought Step Verification at Scale: We introduce a scalable execution-grounded synthesis pipeline that achieves rationale-step verification.
  2. (2)
    Concept-First Curriculum Synthesis: A curriculum-driven synthesis that provides control over problem complexity and diversity, enabling systematic coverage from basic constructs to advanced algorithms.
  3. (3)
    Comprehensive Empirical Validation: Through systematic validations, we demonstrate that execution-groundedness and bi-directional nature of CoTs directly determine model performance.
  4. (4)
    Open-Source Pipeline: Complete synthesis infrastructure including concept extraction, hierarchical generation, Dual Agreement verification, and trace-to-rationale translation to facilitate reproducible research in verified reasoning for code.
The remainder of this paper is organized as follows. We first review related work in code reasoning and synthetic data generation. Next, we detail our three-stage data synthesis pipeline, followed by our experimental design and a thorough analysis of the results. We conclude with a discussion of our findings and future research directions.
2. Related Work
Our work builds upon several interconnected research areas in code reasoning, execution-based program understanding, and synthetic data generation. Chain-of-Thought (CoT) prompting (wei2022chain) has proven effective for improving reasoning, but its extension to code faces unique verification challenges. While methods like Program-of-Thought (PoT) (chen2022program) generate executable snippets, they lack systematic verification of the reasoning process itself. Building on this foundation, CodeI/O (li2025codeio) takes a significant step forward by transforming code patterns into natural language CoT rationales through input/output prediction tasks. By training models to predict inputs and outputs given code, CodeI/O exposes models to universal reasoning primitives while decoupling reasoning from syntax. However, CodeI/O focuses on prediction tasks rather than generating complete reasoning datasets from execution traces.
Verification Approaches in Code Generation. The term ”verification” in code generation literature often refers to different concepts. We identify three distinct approaches:
  1. (1)
    Outcome Verification validates only the final program output against expected results. This is the most common approach, used by benchmarks like HumanEval (chen2021evaluating) and methods like SemCoder (ding2024semcodertrainingcodelanguage) and rStar-Coder (liu2025rstarcoderscalingcompetitivecode). While effective for validating solutions, it offers no guarantee about the logical integrity of intermediate reasoning steps.
  2. (2)
    Formal Verification uses mathematical proofs (e.g., symbolic execution, SMT solvers) to guarantee program correctness. While sound, this approach is computationally heavyweight and impractical for large-scale training data generation.
  3. (3)
    Rationale-Step Verification (our contribution) validates each reasoning step in the CoT chain. By generating rationales directly from a program’s execution trace—a factual record capturing variable values, state transitions, and control flow—we verify that each reasoning step correctly describes runtime states and semantic operations. This verifies the logical integrity of the entire reasoning process, not just the outcome.
Recent work like TRACED (ding2024traced) and Execution Tuning (E.T.) (armengol2025execution) has explored using execution traces for pre-training, but these methods do not generate natural language rationales with step-by-step verification. Our work is the first to systematically achieve rationale-step verification at scale by translating execution traces into verifiable natural language CoT. Conceptually closest to our trace translation, Jung et al. generate natural language CoT from execution traces but focus on single-direction reasoning rather than a systematic bi-directional pipeline.
Comparison to SemCoder and rStar-Coder. SemCoder and rStar-Coder also generate synthetic CoT data, but our methodology differs in two aspects. First, while SemCoder generates rationales via an LLM’s explanation and verifies only output, we perform a direct translation of execution traces, eliminating logical hallucinations. Second, unlike SemCoder’s code-first approach, our concept-first approach provides fine-grained control over complexity and diversity by synthesizing problems from abstract technical concepts. Similarly, rStar-Coder excels at verifiably correct solutions, while we focus on creating verifiably correct reasoning processes.
Bi-Directional Reasoning. The principle of bi-directional reasoning has been explored in code by CodeI/O (li2025codeio) through I/O prediction tasks, and in other domains by FOBAR (jiang2024forward) and REVTHINK (chen2024reverse). However, these approaches focus on prediction tasks rather than generating complete trace-grounded rationales. Other verification techniques like Self-Verification (weng2023large) and RCoT (xue2023rcot) rely on complex checking procedures, while our deterministic trace-based method provides guaranteed correctness.
As summarized in Table 1, no existing work provides a systematic pipeline for generating bi-directional CoT datasets where every reasoning step is grounded in verifiable execution. Our work addresses this gap, providing a foundation for training robust code reasoning capabilities.
3. Data Synthesis Pipeline
Our data synthesis pipeline (Figure 2) is a multi-stage process designed to generate high-fidelity, bi-directional CoT data where every reasoning step is anchored in ground-truth program execution.
3.1. Stage A: Concept Sourcing and Curriculum-Driven Synthesis
@@
Rather than generating code from simple prompts, our pipeline begins by building a curriculum of programming concepts derived from high-quality sources. This curriculum-driven approach ensures the resulting problems are complex, diverse, and grounded in established knowledge.
3.1.1. Document Processing and Concept Extraction
Our pipeline processes a diverse corpus of permissively-licensed technical literature, including books from the StarCoder2-documentation dataset (starcoder2-doc) and curated programming resources spanning basic to advanced topics. Rather than relying on raw PDF text extraction, which produces noisy output, we employ Docling (Docling), a document understanding framework that renders PDFs into clean, structured markdown. This preprocessing preserves semantic structure (headings, code blocks, lists), removes pagination artifacts, and maintains proper formatting. The cleaned text is chunked into 4000-character segments with sliding overlap to prevent concept fragmentation.
3.1.2. Hybrid Concept Identification
For each text chunk, we employ a three-stage hybrid extraction strategy balancing recall with precision:
Stage 1: Statistical Keyword Extraction. We use spaCy with PyTextRank, a graph-based ranking algorithm, to identify candidate concepts. PyTextRank applies the TextRank algorithm (adapted from PageRank) to the chunk’s lemma graph, where nodes represent terms and edges represent co-occurrence relationships. This produces a ranked list of phrases based on their centrality. However, this captures non-technical terms (e.g., page numbers, dates) and misses implicit concepts discussed narratively.
Stage 2: LLM-Based Filtering and Augmentation. To address these limitations, we prompt an LLM teacher model (Qwen2.5-Coder-7B-Instruct) with a filtering prompt that: (1) removes book metadata and non-Python-specific terms, (2) completes incomplete phrases, and (3) identifies implicit concepts present in text but not captured statistically. This hybrid approach yields approximately 15,000 initial concepts.
Why This Hybrid Approach? Pure NLP methods lack semantic understanding and generate noise; pure LLM extraction is expensive and may hallucinate concepts not in text. Our hybrid leverages the efficiency of statistical methods for recall and LLM semantic reasoning for precision.
3.1.3. Rigorous Concept Deduplication and Quality Scoring
The initial 15,000 concepts contain substantial redundancy and quality variance. We apply multi-stage quality control:
Deduplication. We normalize concepts (lowercasing, lemmatizing) and perform string similarity matching to cluster near-duplicates, selecting the most complete variant as canonical.
LLM-Based Quality Scoring. For each deduplicated concept, we prompt the teacher model to score along two dimensions: (1) Difficulty (1-5): algorithmic complexity required for implementation, and (2) Relevance (1-5): ability to inspire diverse, non-trivial problems. We retain only concepts scoring $\geq 3$ on both dimensions, yielding approximately 8,000 high-quality seed concepts.
Why Difficulty Filtering Matters. Preliminary experiments revealed that naively sampling produces a distribution heavily skewed toward trivial problems. This is problematic because: (1) trivial problems do not challenge multi-step reasoning, (2) resulting CoT rationales are too short for meaningful supervision, and (3) downstream verification struggles with trivial problems where nearly all candidates pass all tests. Difficulty filtering biases the curriculum toward problems requiring algorithmic thinking.

### v3

# Think Like You Execute: Verifiable Chain of Thought from Program Traces

@@
Abstract
[…]
Think Like You Execute: Verifiable Chain of Thought from Program Traces
Shailja Thakur1, Vaibhav Saxena1, Rohan Kulkarni1, Shivdeep Singh1, Parameswaran Selvam1, Hiroshi Kanayama2, Hima Patel1 1IBM Research, India 2IBM Research, Japan
1 Introduction
Recent advances in LLMs have enabled code assistants to generate, explain, and debug code Chen et al. (2021); Austin et al. (2021), but they struggle when tasks require reasoning about what a program actually does Ding et al. (2023)—tracking how variables change during debugging, following execution paths to comprehend logic, or confirming that a refactored program preserves its original behavior. When asked to explain code execution, these models often produce plausible but incorrect accounts Turpin et al. (2023).
To improve reasoning capabilities, researchers have turned to Chain-of-Thought (CoT) training Wei et al. (2023a). However, current CoT datasets for code are generated without grounding in actual program execution, producing rationales that sound correct but are untethered to runtime behavior. Figure 1 illustrates how LLMs hallucinate incorrect variable states and control flow despite having access to the code—for instance, comparing “arr[2] with arr[1]” instead of “arr[3]”, and claiming peak direction as “right” instead of “left”.
Code LLM-Generated CoT Execution Trace Trace-Grounded & Verified CoT (Ours)
def find_peak(arr):
 left,right=0,len(arr)-1
while left<right:
 mid=(left+right)//2
if arr[mid]<arr[mid+1]:
  left=mid+1
else: right=mid
return left
# find_peak([1,3,5,4,2])
[…]
✗ Hallucinates
[…]
Return: 2
Starting with array [1,3,5,4,2], we initialize left=0 and right=4. First iteration finds midpoint at index 2 (value 5). Comparing with arr[3]=4 shows we’re past the peak, so we set right=2. Second iteration has mid=1 (value 3). Comparing with arr[2]=5 shows the peak is ahead, so left moves to 2. Boundaries converge at index 2, which holds our peak value of 5. ✓ Verified step-by-step
Figure 1: Hallucinated vs trace-grounded CoT. LLM hallucinates (arr[1] vs arr[3], right vs left). Our approach narrates execution traces with color correspondence (yellow=Init, blue=Loop, green=Return) proving verification.
Prior work has tackled parts of this problem. Some translate execution traces into rationales Ding et al. (2023); Armengol-Estapé et al. (2025) but do not systematically verify the generated reasoning against the trace, while others check final outputs Li et al. (2025); Ding et al. (2024); Liu et al. (2025) but skip intermediate steps. Most also handle only forward reasoning (predicting outputs from inputs) and overlook backward reasoning (inferring inputs from outputs) Li et al. (2025); Jiang et al. (2024). We close these gaps with an execution-grounded CoT synthesis pipeline that translates traces into natural language, cross-checks each rationale against the trace for (1) variable values, (2) state transitions, and (3) control flow, and generates bi-directional rationales for both forward and backward reasoning.
@@ 1 Introduction
Figure 2: Overview of our three-stage data synthesis pipeline. Stage A: Generates candidate concepts, code, and tests from documents. Stage B: Uses Dual Agreement algorithm to identify highest-quality solution-test pairs. Stage C: Produces bi-directional training samples with trace-grounded & verified rationales.
@@
Table 1: Comparison of features in related works. Our work is the first to systematically combine execution-grounded verification with bi-directional CoT generation at scale. Abbreviations: NL=Natural Language.
Work Verifies NL Bidirectional Verifies Scalable
CoT CoT CoT Output Data
REVTHINK Chen et al. (2025)
TRACED Ding et al. (2023)
CodeI/O Li et al. (2025)
rStar-Coder Liu et al. (2025)
Jung et al. Jung et al. (2025)
SemCoder Ding et al. (2024)
@@
Our contributions are:
  1. 1.
    An end-to-end pipeline that generates execution-grounded CoT rationales and verifies each against the trace, demonstrated at scale with 54,000 examples.
  2. 2.
    Systematic bi-directional CoT generation (forward and backward reasoning) enabling comprehensive code understanding.
  3. 3.
    Empirical evidence that verified data quality outperforms quantity, and that execution grounding drives the largest gains in reasoning.
  4. 4.
    Complete synthesis pipeline available as open-source contribution to support reproducibility.
2 Related Work
Our work builds upon code reasoning, execution-based program understanding, and synthetic data generation. Chain-of-Thought prompting Wei et al. (2023a) and Program-of-Thought Chen et al. (2023a) lack systematic verification of reasoning steps. CodeI/O Li et al. (2025) generates CoT rationales but does not verify intermediate reasoning against execution traces.
Literature typically validates only final program outputs (Outcome Verification), offering no guarantee about intermediate reasoning. By verification in our work, we mean: validating each reasoning step against ground-truth execution traces—records of variable values, state transitions, and control flow. Methods like SemCoder Ding et al. (2024) and rStar-Coder Liu et al. (2025) employ outcome verification but lack step-by-step reasoning validation. Recent work like TRACED Ding et al. (2023) and Execution Tuning Armengol-Estapé et al. (2025) use execution traces for pre-training but do not generate natural language CoT with verification. Jung et al. Jung et al. (2025) generate CoT from traces using a teacher model but focus on single-direction reasoning and do not verify generated rationales against the trace. REVTHINK Chen et al. (2025) explores bidirectional reasoning but without execution grounding or verification. Our setting is also distinct from inference-time execution-feedback methods Chen et al. (2023b); Shinn et al. (2023), which improve test-time behavior rather than grounding training data in observed execution. We address these gaps by verifying intermediate reasoning against execution, supporting both reasoning directions, and generating trace-grounded CoT at scale.
3 Data Synthesis Pipeline
As shown in Figure 2, our three-stage pipeline synthesizes code from diverse programming concepts, filtering high-quality code through execution-based consensus, and produce verified bi-directional CoT rationales at scale.
3.1 Stage A: Concept Sourcing and Synthesis
@@
Rather than generating code from simple prompts, our pipeline builds a curriculum of programming concepts from high-quality sources.
Concept Extraction. To ensure broad coverage, we extract programming concepts from permissively-licensed sources: StarCoder2-documentation StarCoder2-Documentation (2024) and curated programming resources EbookFoundation (2024). Using a hybrid approach combining statistical NLP and LLM-based filtering, we retain concepts spanning data structures, algorithms, string manipulation, numerical computation, concurrency, and object-oriented patterns, balanced across difficulty levels from basic to advanced. Quality control via deduplication and scoring yields approximately 8,000 seed concepts.
Problem Synthesis. For each concept, we synthesize complete problem artifacts through a five-step pipeline: (1) generate diverse natural language instructions across problem domains and difficulty levels; (2) produce formal signature skeletons to maintain consistency; (3) generate candidate solutions conforming to agreed interfaces; (4) identify test scenarios including edge cases; (5) generate unit tests with structural constraints for downstream parsing.
Signatures from (2) enforce structural constraints on generated code (Figure 4), enabling function calls to be placed directly inside test assertions (Figure 5). This consistent, parseable structure allows clean I/O and metadata extraction for downstream trace generation and verification.
3.2 Stage B: Consensus-Based Code Selection
Stage A generates candidate solutions and tests that may be incorrect or may have wrong expected outputs due to LLM generation errors. To identify high-quality code-test pairs, we adapt Dual Agreement Chen et al. (2022), which leverages a key insight: if many independently generated solutions all pass the same independently generated tests, it is statistically unlikely that both the solutions and tests are incorrect.
We execute all $m\times n$ solution-test pairs in a sandboxed environment to construct a binary pass/fail matrix $M\in\{0,1\}^{m\times n}$, where $M[i,j]=1$ if solution $i$ passes test $j$, and 0 otherwise. We then partition solutions into clusters where all solutions within a cluster $C_{i}$ have identical pass/fail patterns across all tests. For each cluster, we assign a quality score:
$\text{Score}(C_{i})=|C_{i}|\times|T_{p}(C_{i})|$
where $|C_{i}|$ is cluster size and $|T_{p}(C_{i})|$ is the number of tests all solutions pass. We select the highest-scoring cluster’s canonical solution (shortest and most readable) and its corresponding test suite for downstream CoT generation. See Appendix 6 for a concrete example showing how 5 solutions and 25 tests are clustered: Cluster 1 with 4 solutions passing 18 tests (score 72.0, selected) versus Cluster 2 with 1 solution passing 5 tests (score 5.0, rejected).
3.3 Stage C: Execution-Grounded CoT Generation & Verification
For each solution-test pair from Stage B, we generate training samples by translating execution traces into natural language CoT rationales, then verifying each against the trace.
Trace Generation. We use pysnooper Rachum et al. (2019) to instrument each executable solution, capturing function entry, each executed line, variable assignments with state transitions, and return values. This produces a ground-truth record of the program’s operational semantics.
Trace Sanitization. Raw pysnooper traces contain formatting artifacts (ANSI codes, timestamps, file paths) unsuitable for model consumption. We apply systematic sanitization via regex-based cleaning (see Appendix 6), producing clean traces while preserving semantically meaningful transitions. This achieves approximately 40% trace length reduction on average.
Question and CoT Generation. We extract ground-truth I/O from sanitized traces and generate paired questions: forward (‘‘Given input X, what does the function return?’’) and backward (‘‘What input would cause output Y?’’). We prompt an LLM to narrate the execution trace for forward CoT and perform deductive reasoning from output state for backward CoT. Figure 3 illustrates forward-only and backward-only examples. Notably, the model translates factual execution rather than generating from scratch.
Verification Through Grounding. Since CoT sentences do not map one-to-one to trace lines, we verify each rationale using a sliding window algorithm. For each sentence, we extract verifiable entities (variable names, values, control flow keywords) using an LLM, then check each entity against a lookahead window in the trace. We also string-match the predicted I/O in the rationale against ground-truth I/O from the test case. Rationales failing either check are discarded (see Appendix for algorithm details and rejection examples).
Dataset Assembly. Verified rationales are assembled into three versions: forward-only and backward-only (Figure 3), and bi-directional (exemplified in Appendix Figure 6). Each sample contains a prompt section (blue tags in examples: instruction, code, and question(s)) and a response section (green regions in examples: trace-grounded CoT(s) and predicted I/O).
Language Scope and Transferability. Our current implementation targets Python because it offers a convenient tracing and instrumentation layer through tools such as pysnooper. However, most of the pipeline is language-agnostic: concept sourcing, code/test synthesis, dual-agreement filtering, trace sanitization, rationale verification, and bi-directional data assembly transfer unchanged across languages. The language-specific component is trace acquisition together with extraction of verifiable entities such as variable values, state transitions, and executed control-flow decisions. Extending the method to another language therefore mainly requires replacing the tracing layer and adapting the parser for language-specific execution artifacts. In a language such as Java, this introduces additional engineering considerations including compilation filtering, maintaining shared code-test signatures, and scalable trace collection, but these changes are localized to the execution interface rather than the overall pipeline design.